OpenStudy (anonymous):

Challenge question: What is the area of a circle, when the radius is 8 feet?

OpenStudy (anonymous):

I know the answer already...

OpenStudy (anonymous):

I just want to see if anyone else can answer it.

OpenStudy (ali2x2):

201.06 ft^2

OpenStudy (astrophysics):

\[A = \pi r ^2\] what am I missing?

OpenStudy (ali2x2):

?

OpenStudy (anonymous):

201.0619298ft^2

OpenStudy (ali2x2):

aw i did it like this

OpenStudy (ali2x2):

A=πr2=π·82≈201.06193ft²

OpenStudy (anonymous):

oops, yes it is

OpenStudy (astrophysics):

\[A = 64 \pi~inches^2\]

OpenStudy (ali2x2):

wait i right?

OpenStudy (ali2x2):

ayyyy lmaooooooo

OpenStudy (anonymous):

everyone give each other a medal for participation

OpenStudy (ali2x2):

k

OpenStudy (ali2x2):

i gave yup a medal

OpenStudy (ali2x2):

so i stand here with no medal, noooooo

Nnesha (nnesha):

LAUGHING OUT LOUD. sorry :|

OpenStudy (astrophysics):

lol

OpenStudy (ali2x2):

ayy lmao

OpenStudy (anonymous):

You have a medal...